OpenStudy (anonymous):

Match the terms with the definitions. 4x^2(3x)(x)= -6x^2(-2x^2)(x^2)= 4x^3(-3x^2)= -3x^4(-2x^2)= 2x^2(3x^3) A.)12x^5 B.)6x^6 C.)-12x^5 D.)6x^5 E.)12x^6

OpenStudy (yamyam70):

some properties used in this process. (x) (x) = x^2 x + x = 2x 2x^2 (2x^3) = 4x^5 (add exponentswhen multiplying) combining like terms, 2x^2 + 2x^2 = 4x^2 2x^2 - 2x^2 = 0

OpenStudy (anonymous):

still lost @yamyam70 I know math what so ever

OpenStudy (yamyam70):

PEMDAS rule :)